OpenStudy (anonymous):

Find the exact value. Sin30degrees x tan60degrees-Cos30degrees

OpenStudy (anonymous):

I'm not positive but I believe sin30 is 1/2, tan60 is sqrt3, and cos30 is sqrt3/2, so when you do that you get sqrt3/2-sqrt3/2=0

OpenStudy (anonymous):

thank you so much ! and what about for 2Cos45xSin45 ?

OpenStudy (anonymous):

yes the answer is 0.

OpenStudy (anonymous):

how would I find the exact value for 2Cos45xSin45 I am having difficulties

OpenStudy (anonymous):

2(sqrt2/2)xsqrt2/2=2(2/4)=1 I think

OpenStudy (anonymous):

but isn't cos 45 = 1/sqrt 2 ?

OpenStudy (anonymous):

nevermind ! thank you so much !

OpenStudy (anonymous):

wait i am confused

OpenStudy (anonymous):

Isn't 2(2/4) = 4/8 ? how is it 1 ?

OpenStudy (anonymous):

no... 2 (2/4) = 2 x 2 /4 = 4 / 4 = 1 otherwise you havent times by 2 because 2/4 = 4/8 :)